Alibaba International Academy is considering changes to its class structure in an effort to increase inviting lecturer productivity. The initial schedule had each inviting lecturer teaching five classes per week, with each class meeting in an hour per day on Wednesday, Thursday and Saturday. Each class consist of 25 students. The new schedule has each inviting lecturer teaching only three classes, but each class meets daily on Monday until Friday for an hour. New classes will contain 55 students.


(i) Calculate the inviting lecturer productivity for the initial schedule.

(ii) Calculate the inviting lecturer productivity for the changed schedule.

(iii) Suppose that each inviting lecturer also is required to have two hours of consultation time each day after his/her class. Is this schedule changed will increase or decline current productivity? What is the percentage?


1
Expert's answer
2021-02-10T06:54:59-0500

i)       5 classes per week

5 hours per day

3 days

In each class 25 students

Number of hours = 5*3

                       =15 hours a week

Productivity = No of hours/ no of students

                       15/25= 0.6

                       Or 60%

 

ii)     3 classes per week

3 hours per day

5 days

In each class 55 students

Number of hours = 3*5

                       =15 hours a week

Productivity = No of hours/ no of students

                       15/55= 0.273

                       Or 27.3%

 

 

iii)   3 classes per week

5 hours per day

5 days

In each class 55 students

Number of hours = 5*5

                       =25 hours a week

Productivity = No of hours/ no of students

                       25/55= 0.455

It will decrease productivity from 60 % (current) to 45.5%, this is due to the increased number of students per class.